Calculate the total rotational kinetic energy of all the molecules in 1 mole of air 25.0ºC.

Grade:11

1 Answers

Navjyot Kalra
8 years ago
To find out total rotational kinetic energy E, substitute R/NA for k in the equation E = kT,
E = kT
= (R/NA) T
For one mole of gas (NA =1), the equation E =(R/NA)T will be,
E = RT
To find out the total rotational kinetic energy of all the molecules, substitute 25.0̊ C for temperature T and 8.31 J/ K for R for one mole of gas in the equation E = RT,
E = RT
= (8.31 J/ K) (25.0̊ C)
= (8.31 J /K)(25+273) K
= (8.31 J/ K)(298 K)
= 2476.38 J
Rounding off to three significant figure, the total rotational kinetic of all the molecule will be 2480 J.